An observer wishes to determine the height of a tower. He takes sights at the top of the tower from A and B, which are 50 feet apart, at the same elevation on a direct line with the tower. The vertical angle at point A is 300 and at point B is 400. What is the height of the tower?
1
Expert's answer
2020-09-18T16:10:17-0400

Let the Height of tower be h,

Distance between point B and tower be x


When the observer is looking from point A

tan30"\degree" ="\frac{h}{50+x}"


"50+x=1.732h"

x=1.732h-50...(1)


Now observer is looking from Point B

tan40"\degree" ="\frac{h}{x}"


h=x"\times" tan40"\degree"

=(1.732h-50)"\times 0.839"

h=1.4533h-41.9549


h-1.4533h=-41.9549

-.04533h=-41.9549

h="\frac{41.9549}{0.4533}"


h=92.55feet

Hence the height of tower is 92.55 feet